# Temporal Smoothness-Aware Rate-Distortion Optimized 4D Gaussian Splatting

### v1

@@ Abstract
Dynamic 4D Gaussian Splatting (4DGS) effectively extends the high-speed rendering capabilities of 3D Gaussian Splatting (3DGS) to represent volumetric videos. However, the large number of Gaussians, substantial temporal redundancies, and especially the absence of an entropy-aware compression framework result in large storage requirements. Consequently, this poses significant challenges for practical deployment, efficient edge-device processing, and data transmission. In this paper, we introduce a novel end-to-end RD-optimized compression framework tailored for 4DGS, aiming to enable flexible, high-fidelity rendering across varied computational platforms. Leveraging Fully Explicit Dynamic Gaussian Splatting (Ex4DGS), one of the state-of-the-art 4DGS methods, as our baseline, we start from the existing 3DGS compression methods for compatibility while effectively addressing additional challenges introduced by the temporal axis. In particular, instead of storing motion trajectories independently per point, we employ a wavelet transform to reflect the real-world smoothness prior, significantly enhancing storage efficiency. This approach yields significantly improved compression ratios and provides a user-controlled balance between compression efficiency and rendering quality. Extensive experiments demonstrate the effectiveness of our method, achieving up to 91$\times$ compression compared to the original Ex4DGS model while maintaining high visual fidelity. These results highlight the applicability of our framework for real-time dynamic scene rendering in diverse scenarios, from resource-constrained edge devices to high-performance environments.
@@ 1 Introduction
  • Our method achieves significant compression ratios (up to 91$\times$) for dynamic scenes while maintaining reasonable visual fidelity and offering flexible control over the rate-distortion trade-off suitable for diverse computing platforms.
@@ 2.1 Neural Rendering
Dynamic scene modeling extends this idea into 4D Gaussian Splatting (4DGS), incorporating time as an additional dimension. Dynamic NeRF methods nerfies ; hypernerf learn deformation fields or multi-space embeddings to handle scene motion, but they remain computationally heavy and slow to render. In contrast, 4DGS 4dgs ; yang2024deformable was introduced as a unified representation for spatiotemporal data, using a set of 4D Gaussians with per-Gaussian motion trajectories or a deformation network to render moving scenes in real time. Overall, Gaussian Splatting has rapidly become a promising alternative to NeRF, offering comparable quality with significantly improved rendering speed, which is crucial for interactive applications.
@@ 2.2 Gaussian Splatting Compression
Extending these techniques to dynamic scenes poses greater memory challenges due to temporal variations. Recent specialized solutions include Light4GS liu2025light4gs , combining spatio-temporal pruning with entropy coding; 4DGS-1K yuan20251000+ , optimizing Gaussian reuse and visibility masks for ultra-fast rendering; MEGA zhang2024mega , replacing spherical harmonics with lightweight predictors for significant compression; and LGS liu2024lgs , pruning motion-insensitive Gaussians and redundant deformation fields. These methods highlight the growing focus on efficient and high-quality rendering of dynamic scenes. QUEEN girish2024queen introduces integer-based quantization enabling entropy coding; however, it does not explicitly optimize entropy within a rate-distortion framework, leaving rate-distortion optimization for dynamic Gaussian Splatting unexplored.
@@ 3.1 Preliminary
$\boldsymbol{\mu}_{s}(t)=\boldsymbol{\mu}_{0}+\frac{t}{L}\boldsymbol{\mu}_{disp}$ (3)
where $\boldsymbol{\mu}_{0}$ is a pivot position and $\boldsymbol{\mu}_{disp}$ is the learned translation. Dynamic Gaussians ($\mathcal{G}_{d}$) use a keyframe approach, storing parameters only at sparse, uniformly sampled keyframes $\mathcal{K}=\{t\mid t=nI,n\in\mathbb{Z},t\in\mathcal{T}\}$. Properties at intermediate times $t$ are interpolated: position $\boldsymbol{\mu}_{d}(t)$ using Cubic Hermite Interpolation (CHip), quaternion rotation $\mathbf{q}_{d}(t)$ using Spherical Linear Interpolation (Slerp), and temporal opacity $\sigma_{t}(t)$ using a custom two-Gaussian mixture model (parameterized by means $a_{s}^{o},a_{f}^{o}$ and variances $b_{s}^{o},b_{f}^{o}$) to handle appearance/disappearance. For dynamic Gaussians in Ex4DGS, scale ($\mathbf{s}_{d}$) and SH coefficients ($\mathbf{c}_{d}$) are typically assumed constant over time.
@@
Gaussian and SH Coefficient Masking. To reduce model complexity, redundant Gaussians are pruned using learnable masks $\phi_{i}$. A soft mask $\phi_{i}^{soft}=\text{sigmoid}(\phi_{i})$ is generated, then binarized to a hard mask $\phi_{i}^{hard}\in\{0,1\}$ using a threshold $\phi_{th}$ and the straight-through estimator (STE): $\phi_{i}^{hard}=\text{sg}(\mathbb{I}(\phi_{i}^{soft}>\phi_{th})-\phi_{i}^{soft})+\phi_{i}^{soft}$. This hard mask is applied to key attributes like scale $s_{i}$ and opacity $\alpha_{i}$. The associated pruning cost, contributing to the rate term $\mathcal{L}_{rate}$, is:
$\mathcal{L}_{GSprune}=\frac{1}{N}\sum_{i}\phi_{i}^{soft}$ (4)
Similarly, redundant SH coefficients are adaptively pruned using learnable masks $\theta_{i}^{(l)}$ per Gaussian $i$ and degree $l\geq 1$. Soft masks $\theta_{i}^{(l)soft}=\text{sigmoid}(\theta_{i}^{(l)})$ are binarized to hard masks $\theta_{i}^{(l)hard}$ via STE using a threshold $\theta_{th}$, too. The hard mask is applied to zero out corresponding coefficients $\hat{c}_{i}^{(l)}=\theta^{(l)hard}c_{i}^{(l)}$. The weighted pruning loss contributing to $\mathcal{L}_{rate}$ is:
$\mathcal{L}_{SHprune}=\frac{1}{N}\sum_{i}\sum_{l=1}^{k}\frac{2l+1}{(k+1)^{2}-1}\theta_{i}^{(l)soft}$ (5)
This masking strategy is uniformly applied to both static and dynamic Gaussians, and experiments confirmed effective pruning across both categories. The results are available in Appendix C.
@@
Opacity. Unlike traditional 3D Gaussian methods, Ex4DGS parameterizes dynamic opacity temporally by defining a scale function with center parameters ($a_{s}^{o},a_{f}^{o}$) and variance parameters ($b_{s}^{o},b_{f}^{o}$), effectively acting as an on/off temporal mask multiplied with a base opacity ($\alpha^{d}$). Given that the base opacity for dynamic points shares similar properties with the static opacity ($\alpha^{s}$), we follow Wang et al. wang2024rdo3dgs and apply scalar quantization to the base opacity. However, the temporal center and variance parameters, being expressed in time units rather than opacity values, exhibit different sensitivities to quantization. Empirical tests indicated that variance parameters ($b_{s}^{o},b_{f}^{o}$) were particularly sensitive to quantization artifacts, hence, we avoid aggressive quantization for these parameters to maintain visual fidelity. Detailed empirical validation of this decision is presented in Section 4.3.
@@ 3.3 Wavelet Compression for Dynamic Positions
To efficiently represent the motion trajectories of dynamic Gaussians ($\boldsymbol{\mu}_{d}(t)\in\mathbb{R}^{T\times 3}$), we employ a single-level Haar wavelet transform along the temporal dimension. Formally, given the dynamic positions $\boldsymbol{\mu}_{d}=[\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{T}]^{\top}$ with $\mathbf{p}_{t}\in\mathbb{R}^{3}$, the wavelet transform is applied as:
@@ 3.4 End-to-End Rate-Distortion Optimization
Following Bellé et al. balle2017end , our framework performs end-to-end rate-distortion optimization. Our framework jointly optimizes all relevant parameters by minimizing a composite loss function. These parameters include the original Ex4DGS parameters, those from our differentiated quantization (such as codebooks and entropy models for ECVQ), and the wavelet-compressed dynamic position representations. This total loss, $\mathcal{L}_{total}$, integrates the standard rate-distortion objective with regularization terms inherent to Ex4DGS:
$\mathcal{L}_{total}=\mathcal{L}_{dist}+\lambda_{R}\mathcal{L}_{rate}+\lambda_{reg}\mathcal{L}_{reg}$ (8)
Here, $\mathcal{L}_{dist}$ represents the reconstruction fidelity loss, typically combining L1 loss and D-SSIM between the rendered image $\mathbf{C}$ and the ground truth image $\mathbf{C}_{GT}$:
$\mathcal{L}_{dist}=(1.0-\lambda_{dssim})\mathcal{L}_{L1}(\mathbf{C},\mathbf{C}_{GT})+\lambda_{dssim}(1.0-\text{SSIM}(\mathbf{C},\mathbf{C}_{GT}))$ (9)
The term $\mathcal{L}_{rate}$ quantifies the estimated bitrate cost of the compressed representation. It aggregates several components aimed at reducing model size. These include entropy-based costs derived from quantized attribute representations (e.g. VQ indices) and potentially auxiliary losses that promote sparsity within the model, such as terms encouraging the pruning of redundant Spherical Harmonics coefficients or entire Gaussian primitives. These components are collectively balanced by the rate weighting factor $\lambda_{R}$. When ECVQ is employed, the objective intrinsically balances the rate cost, and the quantization distortion for the respective attributes during codeword selection and codebook learning. Furthermore, $\mathcal{L}_{reg}$ incorporates regularization terms adapted from Ex4DGS ex4dgs to encourage plausible scene dynamics. This includes terms penalizing excessive linear displacement for static points and terms ensuring temporal smoothness in the motion and rotation trajectories defined by the dynamic keyframes.
By optimizing this comprehensive loss function end-to-end, our framework learns the compression mechanisms (quantization, pruning, wavelet processing via its impact on $\mathcal{L}_{dist}$ and $\mathcal{L}_{rate}$) concurrently with the scene representation itself. This joint process directly balances the bitrate costs against the final rendering quality and dynamic consistency, guided by the various lambda hyperparameters.
@@ 4.1 Experimental Settings
Metrics. As in the previous works, we report standard metrics including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) lpips for quantitative evaluation. We also measure the storage size for each compression level to evaluate if the proposed method achieves reasonable performance across different compression strengths. By varying the compression parameters, we generate rate-distortion (RD) curves to comprehensively analyze the trade-off between storage size (rate) and visual quality (distortion). Additionally, we report frames-per-second (FPS) to measure the rendering speed.
Implementation Details. We first train the base model using Ex4DGS ex4dgs until densification terminates. Subsequently, we apply our proposed RD-optimized method, incorporating Gaussian point pruning, adaptive spherical harmonics (SH) pruning, entropy-constrained vector quantization (ECVQ), and the novel pointwise wavelet transform. To systematically analyze performance at varying compression strengths, we define six compression levels (Levels 1–6), adjusting pruning hyperparameters as follows: for Gaussian point pruning, $\lambda_{GSprune}$ is set to [0.05, 0.02, 0.01, 0.005, 0.002, 0.0005]; for SH coefficient pruning, $\lambda_{SHprune}$ is set to [0.5, 0.2, 0.1, 0.05, 0.02, 0.005], respectively. All experiments are conducted on an NVIDIA RTX 3090 GPU.
Figure 1: Average RD performance comparison between our method (Levels 1-6) and the Ex4DGS ex4dgs on the Neural 3D Video (N3V) dataset. Per-scene details are provided in Appendix A.
Figure 2: Average RD performance comparison between our method (Levels 1-6) and the Ex4DGS ex4dgs on the Technicolor dataset. Per-scene details are provided in Appendix A.
@@ 4.2 Rate-Distortion Performance
In this section, we analyze the end-to-end rate-distortion (RD) optimized compression performance of the proposed method. Our primary goal is to leverage the spatio-temporal characteristics of videos to achieve compression efficiency beyond the limits of existing methods, while offering flexible control over the rate-distortion trade-off. This goes beyond simple masking and quantization by incorporating video-specific properties. To achieve different operating points on the RD curve, we generate compressed Gaussians of varying sizes by adjusting the strength of Spherical Harmonics (SH) and Gaussian point pruning and quantization, similar to RD3DGS wang2024rdo3dgs . This is primarily controlled by adjusting the weight $\lambda$ for $\mathcal{L}_{GSprune}$ and $\mathcal{L}_{SHprune}$. Figures 1 and 2 present the rate-distortion curves (PSNR, SSIM, LPIPS vs. Model size in MB) for the N3V dataset li2022n3v and the Technicolor dataset sabater2017techni , respectively. Due to the space constraint, we provide the per-scene results in Table 4. Furthermore, Figures 5 and 4 provide qualitative
@@
comparisons, showcasing the visual fidelity of our method at key compression levels against original Ex4DGS ex4dgs . The results demonstrate that our proposed approach, incorporating techniques like differentiated quantization and the wavelet transform, achieves significant model size reduction (e.g., up to 91$\times$ compression) compared to Ex4DGS while maintaining reasonable visual quality. This validates the effectiveness of our method in providing a flexible trade-off between compression rate and rendering quality suitable for diverse computational requirements.

### v2

@@ Abstract
[…]
@@ 1 Introduction
[…]
@@ 2.1 Neural Rendering
Dynamic scene modeling extends this idea into 4D Gaussian Splatting (4DGS), incorporating time as an additional dimension. Dynamic NeRF methods nerfies ; hypernerf learn deformation fields or multi-space embeddings to handle scene motion, while others utilize representations like temporal tri-planes for compression wu2024tetrirf , but they remain computationally heavy and slow to render. In contrast, 4DGS 4dgs ; yang2024deformable was introduced as a unified representation for spatiotemporal data, using a set of 4D Gaussians with per-Gaussian motion trajectories or a deformation network to render moving scenes in real time. Overall, Gaussian Splatting has rapidly become a promising alternative to NeRF, offering comparable quality with significantly improved rendering speed, which is crucial for interactive applications.
@@ 2.2 Gaussian Splatting Compression
Extending these techniques to dynamic scenes poses greater memory challenges due to temporal variations. Recent specialized solutions include Light4GS liu2025light4gs , combining spatio-temporal pruning with entropy coding; 4DGS-1K yuan20251000+ , optimizing Gaussian reuse and visibility masks for ultra-fast rendering; MEGA zhang2024mega , replacing spherical harmonics with lightweight predictors for significant compression; and LGS liu2024lgs , which prunes Gaussians and deformation fields within a framework highly specialized for surgical scenes using stereo vision. These methods highlight the growing focus on efficient and high-quality rendering of dynamic scenes. QUEEN girish2024queen introduces integer-based quantization enabling entropy coding; however, it does not explicitly optimize entropy within a rate-distortion framework, leaving rate-distortion optimization for dynamic Gaussian Splatting unexplored. Furthermore, QUEEN’s primary objective of per-frame compression for streaming applications fundamentally differs from our goal of compressing the entire dynamic scene model for efficient storage.
@@ 3.1 Preliminary
$\boldsymbol{\mu}_{s}(t)=\boldsymbol{\mu}_{0}+\frac{t}{L}\boldsymbol{\mu}_{\text{disp}}$ (3)
where $\boldsymbol{\mu}_{0}$ is a pivot position and $\boldsymbol{\mu}_{\text{disp}}$ is the learned translation. Dynamic Gaussians ($\mathcal{G}_{d}$) use a keyframe approach, storing parameters only at sparse, uniformly sampled keyframes $\mathcal{K}=\{t\mid t=nI,n\in\mathbb{Z},t\in\mathcal{T}\}$, where $\mathcal{T}$ represents the set of all timestamps in the entire video sequence. Properties at intermediate times $t$ are interpolated: position $\boldsymbol{\mu}_{d}(t)$ using Cubic Hermite Interpolation (CHip), quaternion rotation $\mathbf{q}_{d}(t)$ using Spherical Linear Interpolation (Slerp), and temporal opacity $\sigma_{t}(t)$ using a custom two-Gaussian mixture model (parameterized by means $a_{s}^{o},a_{f}^{o}$ and variances $b_{s}^{o},b_{f}^{o}$) to handle appearance/disappearance. For dynamic Gaussians in Ex4DGS, scale ($\mathbf{s}_{d}$) and SH coefficients ($\mathbf{c}_{d}$) are typically assumed constant over time.
@@
Gaussian and SH Coefficient Masking. To reduce model complexity, redundant Gaussians are pruned using learnable masks $\phi_{i}$. A soft mask $\phi_{i}^{\text{soft}}=\text{sigmoid}(\phi_{i})$ is generated, then binarized to a hard mask $\phi_{i}^{\text{hard}}\in\{0,1\}$ using a threshold $\phi_{\text{thres}}$ and the straight-through estimator (STE): $\phi_{i}^{\text{hard}}=\text{sg}(\mathbb{I}(\phi_{i}^{\text{soft}}>\phi_{\text{thres}})-\phi_{i}^{\text{soft}})+\phi_{i}^{\text{soft}}$. Here, $\text{sg}(\cdot)$ denotes the stop-gradient operation, and $\mathbb{I}(\cdot)$ is the indicator function. This hard mask is applied to key attributes like scale $s_{i}$ and opacity $\alpha_{i}$. The associated pruning cost, contributing to the rate term $\mathcal{L}_{\text{rate}}$ (the full formulation of which is detailed in Appendix˜A), is:
$\mathcal{L}_{\text{GSprune}}=\frac{1}{N}\sum_{i}\phi_{i}^{\text{soft}}$ (4)
Similarly, redundant SH coefficients are adaptively pruned using learnable masks $\theta_{i}^{(l)}$ per Gaussian $i$ and degree $l\geq 1$. Soft masks $\theta_{i}^{(l)\text{soft}}=\text{sigmoid}(\theta_{i}^{(l)})$ are binarized to hard masks $\theta_{i}^{(l)\text{hard}}$ via STE using a threshold $\theta_{th}$, too. The hard mask is applied to zero out corresponding coefficients $\hat{c}_{i}^{(l)}=\theta^{(l)\text{hard}}c_{i}^{(l)}$. The weighted pruning loss contributing to $\mathcal{L}_{\text{rate}}$ is:
$\mathcal{L}_{\text{SHprune}}=\frac{1}{N}\sum_{i}\sum_{l=1}^{k}\frac{2l+1}{(k+1)^{2}-1}\theta_{i}^{(l)\text{soft}},$ (5)
where $k$ is the maximum degree of the SHs, typically set to 3. This masking strategy is uniformly applied to both static and dynamic Gaussians, and experiments confirmed effective pruning across both categories. The results are available in Appendix˜D.
@@
Opacity. Unlike traditional 3D Gaussian methods, Ex4DGS parameterizes dynamic opacity temporally by defining a scale function with center parameters ($a_{s}^{o},a_{f}^{o}$) and variance parameters ($b_{s}^{o},b_{f}^{o}$), effectively acting as an on/off temporal mask multiplied with a base opacity ($\alpha^{d}$). Given that the base opacity for dynamic points shares similar properties with the static opacity ($\alpha^{s}$), we follow Wang et al. wang2024rdo3dgs and apply scalar quantization to the base opacity. However, the temporal center and variance parameters, being expressed in time units rather than opacity values, exhibit different sensitivities to quantization. Empirical tests indicated that variance parameters ($b_{s}^{o},b_{f}^{o}$) were particularly sensitive to quantization artifacts, hence, we avoid aggressive quantization for these parameters to maintain visual fidelity. Detailed empirical validation of this decision is presented in Section˜4.3.
@@ 3.3 Wavelet Compression for Dynamic Positions
To efficiently represent the motion trajectories of dynamic Gaussians ($\boldsymbol{\mu}_{d}(t)\in\mathbb{R}^{T\times 3}$), we employ a single-level Haar wavelet transform along the temporal dimension. The wavelet transform is specifically chosen for its excellent time-frequency localization, enabling an efficient representation of motion trajectories that are generally smooth but may contain localized, high-frequency details (e.g., sharp turns or accelerations). Formally, given the dynamic positions $\boldsymbol{\mu}_{d}=[\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{T}]^{\top}$ with $\mathbf{p}_{t}\in\mathbb{R}^{3}$, the wavelet transform is applied as:
@@ 3.4 End-to-End Rate-Distortion Optimization
Following Bellé et al. balle2017end , our framework performs end-to-end rate-distortion optimization. Our framework jointly optimizes all relevant parameters by minimizing a composite loss function. These parameters include the original Ex4DGS parameters, those from our differentiated quantization (such as codebooks and entropy models for ECVQ), and the wavelet-compressed dynamic position representations. This total loss, $\mathcal{L}_{\text{total}}$, integrates the standard rate-distortion objective with regularization terms inherent to Ex4DGS:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{dist}}+\lambda_{\text{R}}\mathcal{L}_{\text{rate}}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}}$ (8)
Here, $\mathcal{L}_{\text{dist}}$ represents the reconstruction fidelity loss, typically combining L1 loss and D-SSIM between the rendered image $\mathbf{C}$ and the ground truth image $\mathbf{C}_{\text{GT}}$:
$\mathcal{L}_{\text{dist}}=(1.0-\lambda_{\text{dssim}})\mathcal{L}_{L1}(\mathbf{C},\mathbf{C}_{\text{GT}})+\lambda_{\text{dssim}}(1.0-\text{SSIM}(\mathbf{C},\mathbf{C}_{\text{GT}}))$ (9)
The term $\mathcal{L}_{\text{rate}}$ quantifies the estimated bitrate cost of the compressed representation. It aggregates several components aimed at reducing model size. These include entropy-based costs derived from quantized attribute representations (e.g. VQ indices) and potentially auxiliary losses that promote sparsity within the model, such as terms encouraging the pruning of redundant Spherical Harmonics coefficients or entire Gaussian primitives. These components are collectively balanced by the rate weighting factor $\lambda_{\text{R}}$. When ECVQ is employed, the objective intrinsically balances the rate cost, and the quantization distortion for the respective attributes during codeword selection and codebook learning. Furthermore, $\mathcal{L}_{\text{reg}}$ incorporates regularization terms adapted from Ex4DGS ex4dgs to encourage plausible scene dynamics. This includes terms penalizing excessive linear displacement for static points and terms ensuring temporal smoothness in the motion and rotation trajectories defined by the dynamic keyframes.
By optimizing this comprehensive loss function end-to-end, our framework learns the compression mechanisms (quantization, pruning, wavelet processing via its impact on $\mathcal{L}_{\text{dist}}$ and $\mathcal{L}_{\text{rate}}$) concurrently with the scene representation itself. This joint process directly balances the bitrate costs against the final rendering quality and dynamic consistency, guided by the various lambda hyperparameters.
@@ 4.1 Experimental Settings
Metrics. As in the previous works, we report standard metrics including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) lpips for quantitative evaluation. We also measure the storage size for each compression level to evaluate if the proposed method achieves reasonable performance across different compression strengths. By varying the compression parameters, we generate rate-distortion (RD) curves to comprehensively analyze the trade-off between visual quality (distortion) and the storage size (rate) of a single model that includes the entire video sequence. Additionally, we report frames-per-second (FPS) to measure the rendering speed.
Implementation Details. We first train the base model using Ex4DGS ex4dgs until densification terminates. Subsequently, we apply our proposed RD-optimized method, incorporating Gaussian point pruning, adaptive spherical harmonics (SH) pruning, entropy-constrained vector quantization (ECVQ), and the novel pointwise wavelet transform. To systematically analyze performance at varying compression strengths, we define six compression levels (Levels 1–6), adjusting pruning hyperparameters as follows: for Gaussian point pruning, $\lambda_{\text{GSprune}}$ is set to [0.05, 0.02, 0.01, 0.005, 0.002, 0.0005]; for SH coefficient pruning, $\lambda_{\text{SHprune}}$ is set to [0.5, 0.2, 0.1, 0.05, 0.02, 0.005], respectively. All experiments are conducted on an NVIDIA RTX 3090 GPU. On average, with this GPU setup, the initial Ex4DGS pre-training takes approximately 1 hour, and our subsequent ECVQ and pruning optimization stage takes another hour, totaling around 2 hours of training time.
Figure 1: Average RD performance comparison between our method (Levels 1-6) and the Ex4DGS ex4dgs on the Neural 3D Video (N3V) dataset. Per-scene details are provided in Appendix˜B.
Figure 2: Average RD performance comparison between our method (Levels 1-6) and the Ex4DGS ex4dgs on the Technicolor dataset. Per-scene details are provided in Appendix˜B.
@@ 4.2 Rate-Distortion Performance
In this section, we analyze the end-to-end rate-distortion (RD) optimized compression performance of the proposed method. Our primary goal is to leverage the spatio-temporal characteristics of videos to achieve compression efficiency beyond the limits of existing methods, while offering flexible control over the rate-distortion trade-off. This goes beyond simple masking and quantization by incorporating video-specific properties. To achieve different operating points on the RD curve, we generate compressed Gaussians of varying sizes by adjusting the strength of Spherical Harmonics (SH) and Gaussian point pruning and quantization, similar to RD3DGS wang2024rdo3dgs . This is primarily controlled by adjusting the weight $\lambda$ for $\mathcal{L}_{\text{GSprune}}$ and $\mathcal{L}_{\text{SHprune}}$. Figures˜1 and 2 present the rate-distortion curves (PSNR, SSIM, LPIPS vs. Model size in MB) for the N3V dataset li2022n3v and the Technicolor dataset sabater2017techni , respectively. Due to the space constraint, we provide the per-scene results in Table˜6. Furthermore, Figures˜5 and 4 provide qualitative
@@
[…]